Q=12000-10p2 find price elasticitu
1
Expert's answer
2019-04-08T10:07:37-0400

Q = 12000 - 10p2, so p2 = 1200 - 0.1Q.

Ed = -b×p/Q, so Ed = - 0.1×p2/Q.